# Enhancing Emotional Generation Capability of Large Language Models via Emotional Chain-of-Thought

### v2

@@ 1 Introduction
To address the issue of emotional generation assessment, we propose an efficient, answer-free evaluation metric, named Emotional Generation Score (EGS). EGS utilizes Goleman’s theory as a criterion and introduces GPT-3.5  [^1] https://openai.com/chatgpt to evaluate LLMs’ responses in multiple dimensions automatically. With the comparisons on four datasets, we confirm that EGS is consistent to human expert ratings. This observation demonstrates the dependability and trustworthiness of the EGS as an evaluation tool for emotional generation.
@@ 3.2 Emotional Chain-of-Thought
+ECoT 35.20 (normal-$\uparrow$ 3.88) 32.42 (normal-$\uparrow$ 5.29) 31.29 (normal-$\uparrow$ 8.65) 36.52 (normal-$\uparrow$ 12.05) 32.90 (normal-$\uparrow$ 7.85)
@@
+ECoT 34.47 (normal-$\uparrow$ 4.18) 32.58 (normal-$\uparrow$ 3.71) 32.25 (normal-$\uparrow$ 5.31) 35.85 (normal-$\uparrow$ 8.54) 32.11 (normal-$\uparrow$ 8.09)
@@
+ECoT 33.33 (normal-$\uparrow$ 1.32) 31.52 (normal-$\uparrow$ 4.61) 32.19 (normal-$\uparrow$ 3.93) 35.56 (normal-$\uparrow$ 6.12) 31.66 (normal-$\uparrow$ 6.33)
@@
+ECoT 36.42 (normal-$\uparrow$ 2.39) 36.02 (normal-$\uparrow$ 4.40) 36.21 (normal-$\uparrow$ 3.54) 36.70 (normal-$\uparrow$ 3.22) 33.48 (normal-$\uparrow$ 4.95)
@@ 3.3 Evaluation Method
+ECoT 32.35 (normal-$\uparrow$ 3.42) 31.98 (normal-$\uparrow$ 2.74)
@@
+ECoT 34.17 (normal-$\uparrow$ 5.25) 35.22 (normal-$\uparrow$ 7.08)
@@
+ECoT 33.95 (normal-$\uparrow$ 3.46) 36.43 (normal-$\uparrow$ 4.81)
@@ 4.1 Datasets and Models
For the textual tasks, we utilized GPT-4  [^2] https://openai.com/research/gpt-4, ChatGLM3-6B (Du et al., 2022), QWen-14B-Chat (Bai et al., 2023a), LLaMA2-13B (Touvron et al., 2023) as baseline models; for the multimodal tasks, we utilized GPT-4V  [^3] https://openai.com/research/gpt-4v-system-card, QWen-VL (Bai et al., 2023b), LLava-1.5-13B (Liu et al., 2023) as baseline models.
@@ 4.2 Reliability of EGS
+ECoT 36.42 (normal-$\uparrow$ 2.39) 36.02 (normal-$\uparrow$ 4.40) 36.21 (normal-$\uparrow$ 3.54) 36.70 (normal-$\uparrow$ 3.22) 33.48 (normal-$\uparrow$ 4.92)
@@ Appendix A Advance Exploration
Prior to presenting our methods, we conducted a comparative experiment to analyse the emotional generation capabilities of existing Language Model Models (LLMs), including GPT-4  [^4] https://openai.com/research/gpt-4, QWen-14B-Chat Bai et al. (2023a), ChatGLM3-6B (Du et al., 2022), LLaMA2-13B (Touvron et al., 2023).
@@ Appendix E Implementation Details
For the ChatGLM3-6B (Du et al., 2022), QWen-14B-Chat (Bai et al., 2023a), LLaMA2-13B (Touvron et al., 2023), QWen-VL (Bai et al., 2023b), and LLava-1.5-13B (Liu et al., 2023), we set the temperature = 0.1. For the GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-1106-preview) and GPT-4V (gpt-4-1106-vision-preview), we call the API provided by OpenAI  [^5] https://platform.openai.com/docs/api-reference. All experiments were implemented on a NVIDIA A100 GPU.

### v3

@@ 1 Introduction
To address the issue of emotional generation assessment, we propose an efficient, answer-free evaluation metric, named Emotional Generation Score (EGS). EGS utilizes Goleman’s theory as a criterion and introduces GPT-3.5 ***https://openai.com/chatgpt to evaluate LLMs’ responses in multiple dimensions automatically. With the comparisons on four datasets, we confirm that EGS is consistent to human expert ratings. This observation demonstrates the dependability and trustworthiness of the EGS as an evaluation tool for emotional generation.
@@ 3.2 Emotional Chain-of-Thought
+ECoT 35.20 ($\uparrow$ 3.88) 32.42 ($\uparrow$ 5.29) 31.29 ($\uparrow$ 8.65) 36.52 ($\uparrow$ 12.05) 32.90 ($\uparrow$ 7.85)
@@
+ECoT 34.47 ($\uparrow$ 4.18) 32.58 ($\uparrow$ 3.71) 32.25 ($\uparrow$ 5.31) 35.85 ($\uparrow$ 8.54) 32.11 ($\uparrow$ 8.09)
@@
+ECoT 33.33 ($\uparrow$ 1.32) 31.52 ($\uparrow$ 4.61) 32.19 ($\uparrow$ 3.93) 35.56 ($\uparrow$ 6.12) 31.66 ($\uparrow$ 6.33)
@@
+ECoT 36.42 ($\uparrow$ 2.39) 36.02 ($\uparrow$ 4.40) 36.21 ($\uparrow$ 3.54) 36.70 ($\uparrow$ 3.22) 33.48 ($\uparrow$ 4.95)
@@ 3.3 Evaluation Method
+ECoT 32.35 ($\uparrow$ 3.42) 31.98 ($\uparrow$ 2.74)
@@
+ECoT 34.17 ($\uparrow$ 5.25) 35.22 ($\uparrow$ 7.08)
@@
+ECoT 33.95 ($\uparrow$ 3.46) 36.43 ($\uparrow$ 4.81)
@@ 4.1 Datasets and Models
For the textual tasks, we utilized GPT-4 https://openai.com/research/gpt-4, ChatGLM3-6B (Du et al., 2022), QWen-14B-Chat (Bai et al., 2023a), LLaMA2-13B (Touvron et al., 2023) as baseline models; for the multimodal tasks, we utilized GPT-4V https://openai.com/research/gpt-4v-system-card, QWen-VL (Bai et al., 2023b), LLava-1.5-13B (Liu et al., 2023) as baseline models.
@@ 4.2 Reliability of EGS
+ECoT 36.42 ($\uparrow$ 2.39) 36.02 ($\uparrow$ 4.40) 36.21 ($\uparrow$ 3.54) 36.70 ($\uparrow$ 3.22) 33.48 ($\uparrow$ 4.92)
@@ Appendix A Advance Exploration
Prior to presenting our methods, we conducted a comparative experiment to analyse the emotional generation capabilities of existing Language Model Models (LLMs), including GPT-4 §§§https://openai.com/research/gpt-4, QWen-14B-Chat Bai et al. (2023a), ChatGLM3-6B (Du et al., 2022), LLaMA2-13B (Touvron et al., 2023).
@@ Appendix E Implementation Details
For the ChatGLM3-6B (Du et al., 2022), QWen-14B-Chat (Bai et al., 2023a), LLaMA2-13B (Touvron et al., 2023), QWen-VL (Bai et al., 2023b), and LLava-1.5-13B (Liu et al., 2023), we set the temperature = 0.1. For the GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-1106-preview) and GPT-4V (gpt-4-1106-vision-preview), we call the API provided by OpenAI https://platform.openai.com/docs/api-reference. All experiments were implemented on a NVIDIA A100 GPU.